OpenStudy (rock_mit182):

Solve the inequality. Express your answer using interval notation. \[\left| 7k-7 \right|-2<6\]

OpenStudy (rock_mit182):

my answer: (-15/7,1/7)

OpenStudy (rock_mit182):

am i wrong ?

OpenStudy (rock_mit182):

cause the closest answer is : (-1/7,15/7) .....?

OpenStudy (rock_mit182):

A)\[(-\frac{ 1 }{ 7 },\frac{ 15 }{ 7 })\] B) \[(-\infty, -1/7)\] C) \[(-\infty, -1/7)U\frac{ 15 }{ 7 }\] D) \[(-\infty,\frac{ 15 }{ 7 })\]

OpenStudy (rock_mit182):

IS IT A? anyone could walk me through ?

jhonyy9 (jhonyy9):

| 7k-7 | -2 < 6 this is your exercise ?

jhonyy9 (jhonyy9):

like a first step add to both sides 2 and will get |7k-7| < 8 yes ?

jimthompson5910 (jim_thompson5910):

`my answer: (-15/7,1/7)` @rock_mit182 you mixed up the signs somehow

OpenStudy (rock_mit182):

yeah, my mystake was add -7 on both sides instead of adding 7 on both sides

OpenStudy (rock_mit182):

anyway thanks to all of you!

jhonyy9 (jhonyy9):

for example you need to know that |x| = x for x >=0 and -x for x < 0

OpenStudy (rock_mit182):

i appreciate your help!

jhonyy9 (jhonyy9):

was my pleasure good luck bye bye